OpenStudy (anonymous):

Solve by graphing: x^2+5x+6=0

OpenStudy (anonymous):

do you know how y=x^2 looks like?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the graph of y=x^2 + 5x + 6 will look exactly like that but translated several units left and down.

OpenStudy (anonymous):

it says to solve by graphing... you want to do it that way? because it would be easier to just factor.

OpenStudy (anonymous):

My teacher wants us to practice graphing because it will be on the test

OpenStudy (anonymous):

ok... you have your graphing calculator?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

A graph can be made from two points simply find two points of the function and draw a line between them

OpenStudy (anonymous):

That's your graph.

OpenStudy (anonymous):

ever done wolfram alpha?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

here is the graph of your quadratic function y=x^2 + 5x + 6... http://www.wolframalpha.com/input/?i=graph+y+%3D+x%5E2+%2B+5x+%2B+6

OpenStudy (anonymous):

the solution to the eqaution you posted is where the graph crosses the x-axis. What x values does it look like the graph crosses?

OpenStudy (anonymous):

you should be able to convert this to vertex form and graph from there

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

sorry im using the glencoe website as well